OpenStudy (anonymous):

WILL MEDAL!! One beam of electrons moves at right angles to a magnetic field. The force on these electrons is 4.9 × 10-14 newtons. A second beam travels at the same speed, but at a 30° angle with the magnetic field. What force is on these electrons? (4.9 × 10-14 newtons) · tan(30°) (4.9 × 10-14 newtons) · sin(30°) (4.9 × 10-14 newtons) · cos(30°) (4.9 × 10-14 newtons) · arctan(30°) (4.9 × 10-14 newtons) · arccos(30°)

OpenStudy (anonymous):

try the 4th one

OpenStudy (muzzack):

there is two situation... 1)F=qvBsin(theta) in this situation F=4.9*10^-14 N theta =90 degree.....so sin(theta)=1 q(magnitude)=1.6*10^-19 C find B

OpenStudy (anonymous):

actually its the second one

OpenStudy (muzzack):

sin30=0.5

OpenStudy (anonymous):

the 4th one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its the second one

OpenStudy (anonymous):

Thanks @Yopelletsunachan1

OpenStudy (anonymous):

4th

OpenStudy (muzzack):

b i agree @Yopelletsunachan1

OpenStudy (anonymous):

your welcome

OpenStudy (kainui):

I would do this all algebraically: \[F_1=qvB \sin \theta _1\]\[F_2=qvB \sin \theta _2\] Since q the charge of the electron, v the speed, and B the magnetic field don't change we can pretty easily divide the first equation from the second equation to see: \[\frac{F_2}{F_1}=\frac{\sin \theta_2}{\sin \theta _1}\] now just multiply F_1 by both sides and plug in theta 1 and theta 2\[F_2=\frac{\sin \theta_2}{\sin \theta _1}F_1=\frac{\sin 30}{\sin 90}(4.9*10^{-14})\] Sort of saves a lot of pain if you can do it this way I think.